OpenStudy (anonymous):

Simplify: 64 = t - 4 A. t = 16 B. t = 68 C. t = 60 D. None of the above @sangya21

OpenStudy (kamibug):

\(\huge\cal\color{royalblue}{Welcome~to~Openstudy!~:)}\)

OpenStudy (igreen):

Add 4 to both sides..what's 64 + 4?

OpenStudy (anonymous):

THANKS @KamiBug

OpenStudy (anonymous):

64 = t - 4 add 4 to both sides 64+4 = t+(4-4) 68 = t

OpenStudy (anonymous):

68 @iGreen

OpenStudy (igreen):

Yes, so that's your answer.